# Capacity Provisioning Motivated Online Non-Convex Optimization Problem with Memory and Switching Cost

### v1

@@ 2.2. $\alpha\leq 1$
Consider an algorithm $\mathcal{A}_{f}$ that chooses $s(t)=n(t)$ for each time slot.
@@ 2.3. $\alpha>1$
since $\sum_{t}{\mathbf{1}}_{\downarrow}(t)\alpha(-s(t)+s(t-1))$ is at most half of the total switching cost out of the total cost (9) for $\mathcal{A}_{LG}$ because of the linear switching cost and work neutrality constraint. Thus,
@@ 2.5. Algorithm
(23) $\lambda_{j}-\left(\frac{t-a_{j}}{w_{j}}\right)\leq\frac{3}{\beta}(\alpha n_{\textsf{alg}}(t))^{1/2}$
for all $t$ and $\beta\geq 1$, where $n_{\textsf{alg}}(t)$ is the number of remaining jobs with alg at time $t$.
@@
The main advantage of the primal-dual scheme is that there is no need to characterize the structure or any other property of the $\mathsf{OPT}$ directly, which is anyway not available for (20). By exploiting the monotonicity property (Definition 4) and the fact that any new job does not preempt an existing job (which is true since all job sizes are identical) for alg, we derive an upper bound on the primal cost and a lower bound on the dual cost with alg, which together with the weak duality implies the competitive ratio bound.
@@
One obvious question at this point is can we solve this simpler case of Problem 20 when all jobs arrive at the same time, optimally. Next, we answer this question in the affirmative if there is no integrality constraint on the number of active servers $s(t)$. Removing any of these two restrictions, especially the integrality constraint, it is hard to solve the problem exactly even when all jobs arrive at the same time.
To solve (20) exactly when all jobs arrive at the same time, consider the following optimization problem. Let there be $n$ unit sized jobs available at time 00, and problem ${\mathbf{P}}_{H}$ is to minimize the flow time plus the quadratic switching cost over a horizon of $H$ time slots, i.e., all $n$ jobs should be completed by $H$ time slots. [^3] Extension to the case when all jobs are of same size not equal to unity is straightforward. Note that (20) is equivalent to $\min_{H}{\mathbf{P}}_{H}$, when all job sizes are unity. Thus if we can solve ${\mathbf{P}}_{H}$ for any $H$ (structurally) we can solve (20) as well. To solve ${\mathbf{P}}_{H}$, let $s(t)$ denote the number of servers active in time slot $t$ equivalently, $s(t)$ denotes the number of jobs completed in time slot $t.$
Note that with ${\mathbf{P}}_{H}$, the total flow time equals $(n-s(1))+(n-s(1)-s(2))+\cdots,$. Thus, ${\mathbf{P}}_{H}$ is to solve
@@ 5. Conclusions
In this paper, we have broadened the horizon of OCO-S problem by incorporating two important aspects: i) the objective function has unbounded memory, i.e., at any time depends on all the actions taken so far, and ii) the objective function is non-convex, for a specific QoS driven cost function motivated by capacity provisioning where the queuing dynamics interact with server switching cost. Prior work on OCO indicated that the regret or the competitive ratio increases linearly in the memory size, however, using the specific structure of the considered problem, we avoided the negative results, and derived algorithms with memory independent competitive ratios.

### v2

@@ 2.2. $\alpha\leq 1$
Consider an algorithm $\mathcal{A}_{f}$ that chooses $s(t)=n(t)$ for each time slot $t$.
@@ 2.3. $\alpha>1$
since $\sum_{t}{\mathbf{1}}_{\downarrow}(t)\alpha(-s(t)+s(t-1))$ is at most half of the total switching cost out of the total cost (9) for $\mathcal{A}_{LG}$ because of the linear switching cost and work neutrality constraint. Thus, (18) gives
@@ 2.5. Algorithm
(23) $\lambda_{j}-\left(\frac{t-a_{j}}{w_{j}}\right)\leq\frac{3}{\beta}(\alpha n(t))^{1/2}$
for all $t$ and $\beta\geq 1$, where $n(t)$ is the number of remaining jobs with alg at time $t$.
@@
The main advantage of the primal-dual scheme is that there is no need to characterize the structure or any other property of the $\mathsf{OPT}$ directly, which is anyway not available for (20). By exploiting the monotonicity property of alg (an algorithm is defined to be monotone if $n^{\mathcal{I}}(t)\leq n^{\mathcal{I}^{\prime}}(t)$, where $n^{\mathcal{I}}(t)$ is the number of remaining jobs at time $t$ with job arrival instance $\mathcal{I}$ and $\mathcal{I}\subseteq\mathcal{I}^{\prime}$) and the fact that any new job does not preempt an existing job (which is true since all job sizes are identical) for alg, we derive an upper bound on the primal cost and a lower bound on the dual cost with alg, which together with the weak duality implies the competitive ratio bound.
@@
Next, we solve the simpler case of Problem 20 when all jobs arrive at the same time, optimally, as long we do not enforce the integrality constraint on the number of active servers $s(t)$.
To solve (20) exactly when all jobs arrive at the same time, consider the following optimization problem. Let there be $n$ unit sized jobs available at time 00, and problem ${\mathbf{P}}_{H}$ is to minimize the flow time plus the quadratic switching cost over a horizon of $H$ time slots, i.e., all $n$ jobs should be completed by $H$ time slots. $\min_{H}{\mathbf{P}}_{H}$, when all job sizes are unity. Thus if we can solve ${\mathbf{P}}_{H}$ for any $H$ (structurally) we can solve (20) as well. To solve ${\mathbf{P}}_{H}$, let $s(t)$ denote the number of servers active in time slot $t$ equivalently, $s(t)$ denotes the number of jobs completed in time slot $t.$
Note that with ${\mathbf{P}}_{H}$, the total flow time equals $(n-s(1))+(n-s(1)-s(2))+\cdots$. Thus, ${\mathbf{P}}_{H}$ is to solve
@@ 5. Conclusions
In this paper, we have broadened the horizon of OCO-S problem by incorporating two important aspects: i) the objective function has unbounded memory, i.e., at any time depends on all the actions taken so far, and ii) the considered objective function motivated by capacity provisioning where the queuing dynamics interact with server switching cost is non-convex. Prior work on OCO indicated that the regret or the competitive ratio increases linearly in the memory size, however, using the specific structure of the considered problem, we avoided the negative results, and derived algorithms with memory independent competitive ratios.